7. Use dual simplex method to solve the following LPP. (10)

Min 1 2 2 3 3

z = x + x + x

Subject to

x1 − x2 + x3 ≥

x1 + x2 + 2x3 ≤ 8

x1 − x3 ≥ 2

x1

, x2

, x3 ≥ .0


1
Expert's answer
2021-05-07T09:55:03-0400

Min "Z = x_1+2x_2+3x_3"


Subject to :

"x_1 − x_2 + x_3 \ge 2"


"x_1 + x_2 + 2x_3 \le 8"

"x1 − x3 \le 2"

"x_1,x_2,x_3 \ge 0"

After introducing slack,surplus, artificial variables

Max "Z = x_1+2x_2+3x_3+0S_1+0S_2+ 0S_3 -MA_1 - MA_2"

Subject to

"x_1 - x_2+ x_3 - S_1+ A_1 = 2"

"x_1+x_2+2x_3+S_2 = 8"

"x_1-x_3-S_3+A_2 = 2"

and all are greater than zero.




Negative number "Z_j - C_j" is -"2M-1" and its column index is 1. So, the entering variable is "x_1" .

Minimum ratio is 2 and its row index is 1. So, the leaving basis variable is "A_1" .

Therefore, the pivot element is 1.

Entering = "x_1" , Departing = A_1 , Key element = 1

"R_1(new) = R_1(old)"

"R_2(new) = R_2(old) - R_1(new)"

"R_3(new) = R_3(old)- R_1(new)"





Negative number "Z_j - C_j" is "-M-3" and its column index is 2. So, the entering variable is "x_2" .

Minimum ratio is 0 and its row index is 3. So, the leaving basis variable is "A_2" .

Therefore, the pivot element is 1.

Entering = "x_2" , Departing = "A_2" , Key element = 1

"R_3(new) = R_3(old)"

"R_1(new) = R_1(old) + R_1(new)"

"R_2(new) = R_2(old)- 2R_3(new)"






Negative number "Z_j - C_j" is -8 and its column index is 3. So, the entering variable is "x_3" .

Minimum ratio is 1.2 and its row index is 2. So, the leaving basis variable is "S_2" .

Therefore, the pivot element is 5.

Entering = "x_3" , Departing = "S_2" , Key element = 5


"R_2(new) = \dfrac{R_2(old)}{5}"

"R_1(new) = R_1(old) + R_2(new)"

"R_3(new) = R_3(old)+2R_2(new)"






Since all "Z_j-C_j \ge 0."

Hence, the optimal solution is arrived with value of variables as:

"x_1 = 3.2"

"x_2 = 2.4"

"x_3 = 1.2"

Max "Z = 11.6"